find the orthogonal trajectories of the family of hyperbolas xy=c (c not= 0 )
1
Expert's answer
2020-09-28T20:07:00-0400
"Solution"


From the given family of curves, we find a differential equation the curves all satisfy,


"xy=C \implies y'=-\frac{C}{x^2}=-\frac{y}{x}"

Letting "f(x,y)=-\frac{y}{x}", we know the orthogonal trajectories are the curves which satisfy a differential equation


"y'=-\frac{1}{f(x,y)}\implies y'=\frac{x}{y}"

Therefore, the orthogonal trajectories are the curves,


"y'=\frac{x}{y} \implies yy'=x \implies \frac12y^2 =\frac12x^2+C\\
\implies y^2-x^2=C"

where "C" is an arbitrary constant.